# A Survey on SAR ship classification using Deep Learning

### v1

@@ IV-D Fine-Tuning
High resolution images from the CosmoSkyMed mission (level 1A HIMAGE single-look complex Slant (SCS) product images in the X-band with single HH or VV polarization) are exploited to test the effectiveness of pre-trained models on a small dataset with three classes. [22] conclude that fine-tuning a VGG16 network can lead to f1-scores of up to 97.8%. This shows the usefulness of DL techniques for datasets with less samples like SAR.
@@
[24] demonstrates that leveraging transfer learning with pre-trained models, combined with effective data augmentation strategies, can significantly improve the classification accuracy even when working with small training datasets, achieving an accuracy of 98% using a deep architecture, DenseNet-121. The input images are sourced from the TerraSAR-X stripmap mode and are manually annotated by interpretation experts using AIS data. There are only three classes: bulk carrier, container ship, and oil tanker. The data augmentation techniques applied are flipping, brightening, sharpening, random cropping, and rotating.
@@
SAR data acquisition involves varying incident angles. [25] examines the effect of these varying angles on model performance. It uses a dataset containing images of three classes taken from Cosmo-SkyMed and GaoFen-3, encompassing incident angles ranging from 20 to 60 degrees. The data were divided into three overlapping subsets: one containing the entire dataset, another containing data with low incident angles (less than 48 degrees), and a third one containing data with high incident angles (greater than 45 degrees) . All three datasets were tested on four pre-trained DL models: VGG-16, VGG-19, ResNet-50, and ResNet-101. The results indicate that incorporating data from all available angles enhances the models’ predictive accuracy.
@@
Data scarcity can be addressed by training the model on domain-specific data, such as optical images of ships, and then applying domain adaptation to SAR data. [44] proposes a novel method, DSAN++, for domain adaptation. This method involves first training on optical images and then transferring the learned knowledge to SAR data. DSAN++ utilizes two ResNet networks for feature extraction. The extracted features undergo feature fusion, followed by the application of a subdomain adaptation network to learn transferable representations. To capture both local and global distributions, shallow and deep ResNet models (ResNet-18 and ResNet-50) are integrated with the convolutional block attention module. The experiments were conducted by training the model on a four-category (cargo, bulk carrier, container ship, and oil tanker) optical remote sensing (ORS) dataset obtained from Google Earth. Subsequently, domain adaptation was performed on two datasets: (1) GF-SAR with three classes (bulk carrier, container ship, and oil tanker) collected from FUSAR and GaoFen, and (2) a three-class dataset (cargo, container ship, and oil tanker) derived from TerraSAR-X. The results demonstrate that DSAN++ leads to improved performance in subdomain adaptation. [52] compared the fine-tuning performance of models pre-trained on domain-specific versus non-domain-specific datasets using CNN’s. The methodology was evaluated on three-class targets (carrier, container, and oil tanker) from VV polarized TerraSAR-X, and four-class targets (cargo, bulk carrier, container, and oil tanker) from FUSARShip. The results suggest that shallow models with feature refinement can achieve better performance when fine-tuned on domain-specific models. The authors investigate the pre-training effects of ImageNet and an ORS dataset with data augmentation techniques (scaling, flipping, and rotating). While ImageNet, due to its large size, yields better performances on deep networks, the ORS dataset enhanced the performance of shallow networks, indicating that shallow models are more effective with less data.
@@ IV-D1 Data Augmentation & GANs
Techniques like data augmentation, when combined with transfer learning, can address data-related problems. [39] demonstrates that using transfer learning with data augmentation and manual features can improve the performance of a DL model. It uses ground range detected SAR data acquired in interferometric wide-swath mode from OpenSARShip. The data consist of two channels, with VH and VV polarizations, stacked together. The study focuses on classifying ships into four categories: bulk carrier, container ship, fishing vessel, and tanker. The authors use five data augmentation techniques (horizontal flipping, vertical flipping, rotation by 90°, rotation by 270°, and Gaussian noise injection) and incorporate 14 hand-crafted features. The effectiveness of these hand-crafted features combined with DL features are compared across 3 machine learning (Radial Basis Function (RBF)-SVM with two stage feature selection, SVM classifier combination, SVM one-vs-rest) and 4 DL models (CNN, AlexNet, VGG19, and ResNet50). Additionally, the authors present a methodology for matching AIS and SAR data and test the performance on real data collected from three different locations. Their results show that the pre-trained ResNet50 model, when fine-tuned on SAR images and hand-crafted features, achieves an accuracy of 82.5%.
@@
One of the possible DL-based solutions to the data scarcity problem in optical images is represented by generative models and can also be applied in SAR images. These models learn data representations and produce images that resemble the originals, effectively increasing the available data. [45] proposes a novel conditional generative adversarial network (GAN) method that generates high-quality SAR images, improving ship classification performance. The method consists of two parts: (1) a residual convolutional block is integrated with the generator to enhance the detail of generated images, and (2) the generated images are merged with real images and sent to a discriminator for two tasks: distinguishing real from fake images and classifying them into three classes. The authors use a three-class dataset (cargo, container, tanker) from high-resolution SAR images of the Chinese C-band, 3-meter resolution GaoFen-3 platform. The results indicate that the GAN is capable of generating high-resolution images, and that these images, coupled with the real images, enhance the classification performance.
@@
[65] proposes a generative model-based domain adaptation framework for ship classification on a 5-class FUSARShip dataset. This framework leverages domain adaptation by transferring knowledge from the optical images of the FGSCR42 dataset. [63] introduces the Attention-Dense CycleGAN (ADCG) to improve recognition by leveraging transfer learning from optical to SAR domains (OPT2SAR). ADCG enhances CycleGAN with a dense connection module (DCM) and a convolutional block attention module (CBAM). Using FUSARShip and the FGSCR optical dataset, ADCG achieves superior performance, evaluated via Fréchet inception distance (FID) and kernel inception distance (KID). It demonstrates a 6% improvement in recognition accuracy for SAR ship classification networks, thus validating its effectiveness in addressing class imbalance and sample scarcity through effective transfer learning.
@@ V-A Datasets
The literature explores two public datasets, FUSARShip and OpenSARShip, and some additional, distinct data sources for SAR ship classification. Their detailed information is provided below.
@@ V-A1 OpenSARShip
The OpenSARship dataset is a publicly available resource designed to support research in ship classification using SAR imagery. This dataset provides a collection of pre-processed SAR image chips specifically focused on ship detection and classification tasks. The current version contains 34528 vessel instances. The data are collected using two products from ESA’s Sentinel-1 repository (single look complex, and ground range detected) in interferometric wide swath mode. The dataset consists of 16 ship classes, containing classes like Cargo, Tanker, Fishing, Tug, Dredging. However, the dataset suffers from class imbalance, with 61% data belonging to Cargo category. Three classes represent 95% of data, while the remaining 13 classes only represent a 5% of the data provided. This imbalance can lead any model to bias toward the majority class. The literature shows that researchers often address this problem by only using the 3 to 5 most populated classes, e.g., cargo, tanker, tug, fishing, and dredging [23, 29, 57, 34, 36, 43, 60], or by utilizing subclasses within the cargo category, e.g., general cargo, container ship, bulk carrier [30, 33, 47]. OpenSARship data incorporates polarization information, VH and VV, which researchers can leverage to create fusion networks [50, 53, 40, 46, 49, 56, 51, 55, 66]. The dataset’s main weaknesses such as low object resolution (20 m/pixel) and class imbalance, limit the performance of a DL model. Nevertheless, if these limitations are addressed, OpenSARship serves as a valuable starting point for researchers to explore their methodologies before acquiring commercial SAR imagery.
@@ V-A2 FUSARShip
The FUSARShip dataset provides a valuable resource for researchers in SAR ship classification. This dataset leverages high-resolution SAR images captured by China’s GaoFen-3 (GF-3) satellite, specifically designed for ocean remote sensing and marine monitoring tasks. FUSARShip is constructed from a collection of image chips extracted from over 126 GF-3 scenes, encompassing 15 primary ship categories, e.g. cargo, tanker, fishing, and 98 subcategories. Each ship instance has a size of $512\times 512512 × 512$ pixels, with a pixel size of 0.5 meters. Similar to OpenSARship, FUSARShip presents a challenge due to class imbalance. Three subcategories constitute approximately an 81% of the data, while the remaining 95 categories collectively only represent a 19% thereof. This significant imbalance can lead to models exhibiting bias towards the overrepresented classes. Consequently, researchers often address this issue by focusing on the 3-5 most populated classes, e.g., cargo, fishing, tanker, bulk carrier [58, 54, 53, 44, 52, 65, 63] as observed in the OpenSARship dataset.
@@ V-A3 Other Data Sources
Large datasets are crucial for training a DL model. However, the imbalance data limitation in the above datasets can hinder model performance. Researchers address this challenge using two techniques. Firstly, they leverage high-quality data from various satellite missions, see Table III, such as TerraSAR-X [21, 26, 24, 44, 52], CosmoSkyMed [22], PALSAR-2 [27], and GaoFen-3 [41, 35, 25, 45]. These missions not only provide high quality data, but also allow researchers to exploit AIS data to strategically select images based on specific locations and timeframes. For instance, if a researcher aims to develop a model for detecting fishing activity during a particular season in a specific region, AIS data can be analyzed to identify relevant satellite images for purchase.
Secondly, due to the cost associated with acquiring high-resolution commercial imagery, researchers often employ transfer learning techniques. This involves training a model on a large, general-purpose dataset (e.g., ImageNet), or domain specific dataset, e.g., MSTAR [75] (SAR images of tanks), optical remote sensing images of ships taken from Google Earth [78] and then on a smaller, task-specific dataset (e.g., FUSARShip). Pre-training on a comprehensive dataset equips the model with the ability to extract generic features from images. Subsequent fine-tuning on the target dataset refines these features for the specific classification task.
@@
- StripMap: $5\times 55 × 5$ m
- IW: $5\times 255 × 25$ m
- EW: $20\times 4020 × 40$ m
@@
- StripMap: $3\times 33 × 3$ m
- Spotlight: $1\times 31 × 3$ m
@@ V-B Data Augmentation
Due to the limited availability of SAR images and the inherent data requirements of DL models, data augmentation techniques have been proposed in the literature to address this challenge. These techniques can improve the performance of DL models for SAR ship classification.
@@ V-B1 Rotation
In the literature, rotation is widely employed to enhance the robustness of DL models for SAR ship classification [39, 49, 56, 54, 34, 21, 53]. It involves artificially rotating the training images by various angles to simulate real-world scenarios where objects might be oriented differently relative to the SAR sensor. A common approach is to rotate images within a range of +4545+45+ 45 to 4545-45- 45 degrees, as this range captures a significant portion of the potential ship orientations. This augmentation strategy helps the model learn features that are invariant to rotation, improving its generalization ability and reducing the risk of overfitting to the specific orientations present in the original training data.
@@ V-B3 Scaling
DL networks typically require a fixed-size input. Image scaling, a data augmentation technique that resizes images to a standard size, addresses this requirement. Scaling can be performed by increasing or decreasing image size, with larger images demanding significantly more computational resources for processing [52, 53]. In SAR ship classification, a common practice is to scale images to a size of 128x128 pixels, balancing accuracy and computational efficiency.
@@ V-B6 Cropping
Naturally, SAR images contain a lot of useless information like background, and the object representations are very small, leading the model to learn non-informative details. Cropping is the process of extracting a specific region of interest containing ships and reducing the useless information like background sea. Cropping is commonly employed in literature to augment data [24, 28, 53, 20].
@@ V-C Handcrafted Features
Whereas augmentation is important to help models learn from more data, it can also lead to overfitting. Moreover, augmenting low-quality images often produces additional low-quality data, which can confuse the model and hinder effective learning. In such scenarios, handcrafted features can provide valuable support by being fused with deep learning features within the model (see Fig. 5).
DL models offer significant advantages by virtue of their inherent feature extraction capabilities, but their performance in SAR ship classification can be hindered by data scarcity and imbalance. To address this challenge, researchers have leveraged handcrafted features. Here are some examples of features identified through our literature search:
@@ V-C1 Histogram of oriented gradients
Histogram of Gradients (HOG, see Fig. 3) features are a powerful tool for extracting informative features from images, particularly those containing well-defined edges and shapes [31, 43, 37]. These features describe the local structure of an image by capturing the orientation and intensity of gradients.

### v2

@@ IV-D Transfer Learning and Domain Adaptation
High resolution images (e.g. from CosmoSkyMed mission, level 1A HIMAGE single-look complex slant (SCS) in the X-band, single HH or VV polarization) can be exploited to test the effectiveness of pre-trained models on datasets that have a limited number of classes (e.g. three). In [22] the authors follow this approach and conclude that fine-tuning a VGG-16 network can lead to F1-scores of up to 97.8%. This shows the usefulness of DL techniques for small-sized datasets (like usual SAR imagery datasets).
@@
In [24], the authors demonstrate that leveraging transfer learning with pre-trained models (based on DenseNet-121 architecture), combined with effective data augmentation strategies, can significantly improve the classification accuracy, even when working with small-sized training datasets. Stripmap input images were sourced from TerraSAR-X mission and have been manually annotated by experts using AIS data. Only three classes are employed (bulk carrier, container ship, and oil tanker), and data augmentation techniques are applied (flipping, brightening, sharpening, random cropping, and rotating). Despite the tiny size of the training dataset, the authors achieved a 98% accuracy.
@@
In [25], the authors examine the effect of varying incident angles on model performance. They use a dataset containing three classes of images taken from Cosmo-SkyMed and GaoFen-3, encompassing incident angles ranging from 20 to 60 degrees. The data were divided into three overlapping subsets: one containing the entire dataset, a second one containing data with low incident angles (less than 48 degrees), and a third one containing data with high incident angles (more than 45 degrees). All three datasets were tested on four pre-trained DL models: VGG-16, VGG-19, ResNet-50, and ResNet-101. The results indicate that incorporating data from all available angles enhances the models’ predictive accuracy.
@@
Data scarcity can be addressed by training the model on domain-specific data, such as optical images of ships, then applying domain adaptation to SAR data. In [47], DSAN++, a novel method for domain adaptation, is proposed. This first involves training on optical images and then transferring the learned knowledge to SAR data. DSAN++ utilizes two ResNet networks for feature extraction. The extracted features undergo feature fusion, followed by the application of a subdomain adaptation network to learn transferable representations. To capture both local and global distributions, shallow and deep ResNet models (ResNet-18 and ResNet-50) are integrated with the convolutional block attention module. The experiments were conducted by training the model on an optical remote sensing (ORS) dataset obtained from Google Earth, with four classes (cargo, bulk carrier, container ship, and oil tanker). Subsequently, domain adaptation was performed on two datasets: (1) GF-SAR with three classes (bulk carrier, container ship, and oil tanker) collected from FUSAR and GaoFen, and (2) a three-class dataset (cargo, container ship, and oil tanker) derived from TerraSAR-X. The results demonstrate that DSAN++ leads to improved performance in subdomain adaptation. In [73], the authors apply domain adaptation from OpenSARship to FuSARShip and vice versa on multiple scales resulting in better performance for both datasets as compared to the one obtained by training and testing on a single dataset. In [55], the authors compare the fine-tuning performance of models pre-trained on domain-specific versus non-domain-specific datasets using CNNs. The methodology was evaluated on three-class targets from VV polarized TerraSAR-X (carrier, container, and oil tanker), and four-class targets from FUSARShip (cargo, bulk carrier, container, and oil tanker). The results suggest that shallow models with feature refinement can achieve better performance when fine-tuned on domain-specific models. The authors investigate the pre-training effects of ImageNet and an ORS dataset with data augmentation techniques (scaling, flipping, and rotating). Whereas ImageNet, due to its large size, yields better performances on deep networks, the ORS dataset enhanced the performance of shallow networks, indicating that shallow models are more effective with less data.
@@ IV-D1 Data Augmentation & GANs
Techniques like data augmentation, when combined with transfer learning, can address data-related problems. In [39], the authors demonstrate that using transfer learning with data augmentation and manual features can improve the performance of a DL model. They use ground range detected SAR data acquired in interferometric wide-swath mode from OpenSARShip. The data consist of two channels, with VH and VV polarizations, stacked together. The study focuses on classifying ships into four categories: bulk carrier, container ship, fishing vessel, and tanker. The authors use five data augmentation techniques (horizontal flipping, vertical flipping, rotation by 90°, rotation by 270°, and Gaussian noise injection) and incorporate 14 handcrafted features. The effectiveness of these handcrafted features combined with DL features are compared across three machine learning and four DL models.The ML models are Radial Basis Function (RBF)-SVM with two stage feature selection, SVM classifier combination, SVM one-vs-rest; the DL models are CNN, AlexNet, VGG-19, and ResNet50. Additionally, the authors present a methodology for matching AIS and SAR data and test the performance on real data collected from three different locations. Their results show that the pretrained ResNet50 model, when fine-tuned on SAR images and handcrafted features, achieves an accuracy of 82.5%.
@@
One of the possible DL-based solutions to the data scarcity problem in optical images is represented by generative models and can also be applied to SAR images. These models learn data representations and produce images that resemble the originals, effectively increasing the available data. In [48], the authors propose a novel conditional generative adversarial network method (GAN) that generates high-quality SAR images, improving ship classification performance. The method consists of two parts: (1) a residual convolutional block is integrated with the generator to enhance the detail of generated images, and (2) the generated images are merged with real images and sent to a discriminator for two tasks: distinguishing real from fake images and classifying them into three classes. The authors use a three-class dataset (cargo, container, tanker) from GaoFen-3. The results indicate that the GAN is capable of generating high-resolution images that, coupled with the real images, enhance the classification performance.
@@
A generative model-based domain adaptation framework for ship classification on a 5-class FUSARShip dataset is proposed in [68]. This framework leverages domain adaptation by transferring knowledge from the optical images of the FGSCR42 dataset. In [66], the Attention-Dense CycleGAN (ADCG) is introduced to improve recognition by leveraging transfer learning from optical to SAR domains (OPT2SAR). ADCG enhances CycleGAN with a dense connection module (DCM) and a convolutional block attention module (CBAM). Using FUSARShip and the FGSCR optical dataset, ADCG achieves superior performance, evaluated via Fréchet inception distance (FID) and kernel inception distance (KID). It demonstrates a 6% improvement in recognition accuracy for SAR ship classification networks, thus validating its effectiveness in addressing class imbalance and sample scarcity through effective transfer learning.
@@ V-A Datasets
The literature explores two public datasets, FUSARShip and OpenSARShip, and some additional, distinct data sources for SAR ship classification. More detailed information is provided below.
@@ V-A1 OpenSARShip
The OpenSARship dataset is a publicly available resource designed to support research in ship classification using SAR imagery. This dataset provides a collection of pre-processed SAR image chips specifically focused on ship detection and classification tasks. The current version contains 34528 vessel instances. The data are collected using two products from ESA’s Sentinel-1 repository (single look complex, and ground range detected) in interferometric wide swath mode. The dataset consists of 16 ship typologies, among which classes like Cargo, Tanker, Fishing, Tug, Dredging can be found. However, the dataset suffers from class imbalance, with 61% data belonging to Cargo category. Three classes represent 95% of data, while the remaining 13 classes only represent a 5% of the data provided. This imbalance can lead any model to bias toward the majority class. The literature shows that researchers often address this problem by only using the 3 to 5 most populated classes, e.g., cargo, tanker, tug, fishing, and dredging [23, 29, 60, 34, 36, 46, 63], or by utilizing subclasses within the cargo category, e.g., general cargo, container ship, bulk carrier [30, 33, 50]. OpenSARship data incorporate polarization information, VH and VV, which can be leveraged to create fusion networks [53, 56, 40, 49, 52, 59, 54, 58, 69]. The dataset’s main weaknesses such as low spatial resolution (20 m/pixel) and class imbalance, limit the performance of any DL model. Nevertheless, if these limitations are addressed, OpenSARship serves as a valuable starting point for researchers to explore their methodologies before acquiring commercial SAR imagery.
@@ V-A2 FUSARShip
FUSARShip provides a valuable resource for researchers in SAR ship classification. This dataset uses high-resolution SAR images captured by China’s GaoFen-3 (GF-3) satellite, specifically designed for ocean remote sensing and marine monitoring tasks. FUSARShip is constructed from a collection of image chips extracted from over 126 GF-3 scenes, encompassing 15 primary ship categories and 98 subcategories. Each ship instance has a size of $512\times 512$ pixels, with a pixel size of 0.5 meters. Similar to OpenSARship, FUSARShip presents a challenge due to class imbalance. Three sub-categories constitute approximately an 81% of the data, while the remaining 95 categories collectively only represent a 19% thereof. This significant imbalance can lead to models exhibiting bias towards the overrepresented classes. Consequently, researchers often address this issue by focusing on the 3-5 most populated classes, e.g., cargo, fishing, tanker, bulk carrier [61, 57, 56, 47, 55, 68, 66] as observed in the OpenSARship dataset.
@@ V-A3 Other Data Sources
Large datasets are crucial for training a DL model. However, the imbalance data limitation in the above datasets can hinder model performance. Researchers address this challenge using two techniques. Firstly, they leverage high-quality data from various satellite missions, such as TerraSAR-X [21, 26, 24, 47, 55], CosmoSkyMed [22], PALSAR-2 [27], and GaoFen-3 [41, 35, 25, 48], see Table III. These missions not only provide high quality data, but also allow researchers to exploit AIS to strategically select images based on specific locations and timeframes. For instance, if a researcher aims to develop a model for detecting fishing activity during a particular season in a specific region, AIS data can be analyzed to identify relevant satellite images for purchase.
Secondly, due to the cost associated with acquiring high-resolution commercial imagery, researchers often employ transfer learning techniques. This involves training a model on a large, general-purpose dataset (e.g., ImageNet), or domain specific dataset, e.g., MSTAR [101] (SAR images of tanks), optical remote sensing images of ships taken from Google Earth [103] and then on a smaller, task-specific dataset (e.g., FUSARShip). Pretraining on a comprehensive dataset equips the model with the ability to extract generic features from images. Subsequent fine-tuning on the target dataset refines these features for the specific classification task.
@@
- StripMap: $5\times 5$ m
- IW: $5\times 25$ m
- EW: $20\times 40$ m
@@
- StripMap: $3\times 3$ m
- Spotlight: $1\times 3$ m
@@ V-B Data Augmentation
Due to the limited availability of SAR images and the inherent data requirements of DL models, data augmentation techniques have been proposed in the literature to address this challenge. These techniques can be employed in an attempt to improve the performance of DL models for SAR ship classification.
@@ V-B1 Rotation
in the literature, rotation is widely employed to enhance the robustness of DL models for SAR ship classification [39, 52, 59, 57, 34, 21, 56]. It involves artificially rotating the training images by various angles to simulate real-world scenarios where objects might be oriented differently relative to the SAR sensor. A common approach is to rotate images within a range of +45+45 to 45-45 degrees, as this range captures a significant portion of the potential ship orientations. This augmentation strategy helps the model learn features that are invariant to rotation, improving its generalization ability and reducing the risk of overfitting to the specific orientations present in the original training data.
@@ V-B3 Scaling
DL networks typically require a fixed-size input. Image scaling, a data augmentation technique that resizes images to a standard size, addresses this requirement. Scaling can be performed by increasing or decreasing image size, with larger images demanding significantly more computational resources for processing [55, 56]. In SAR ship classification, a common practice is to scale images to a size of $128\times 128$ pixels, balancing accuracy and computational efficiency.
@@ V-B6 Cropping
maritime SAR imagery contains a lot of information that is useless for ship classification purposes. In fact, the representation of the ship target is usually way smaller than the surrounding background, and this leads the model to learn non-informative details. As a possible countermeasure to reduce useless information, cropping, i.e. the process of extracting a region of interest containing the ship target only, is adopted. Cropping is commonly employed in literature as a data augmentation technique [24, 28, 56, 20].
@@ V-C Handcrafted Features
While data augmentation is often used to increase training diversity, it can also lead to overfitting, especially when applied to low-quality SAR images. In such cases, augmentation may generate additional noisy samples that confuse the model rather than improve it. Handcrafted features provide an alternative source of information that can be fused with DL features to strengthen classification robustness (see Fig. 9).
DL models excel at automatically extracting features, yet their performance in SAR ship classification is constrained by data scarcity, imbalance, and strong speckle noise. To address these issues, researchers have integrated handcrafted features that embed prior knowledge about ship structure and radar scattering. Below, we review the most commonly used handcrafted features in the literature.
@@ V-C1 Histogram of oriented gradients
Histogram of Oriented Gradients (HOG, see Fig. 7) captures edge orientations and local gradient structures [31, 46, 37]. In SAR images, where ship targets typically exhibit strong geometric edges, HOG can highlight hull and superstructure outlines. However, speckle noise can degrade gradient estimation, making HOG alone unreliable. Recent works therefore combine HOG with DL features, where it contributes complementary edge information that can guide the network’s learning of ship contours.